The Ekonland economy is divided into two segments: C and NC, where C is covered by minimum wage, while NC is not. Suppose Ekonland has a fixed number of workers, 1000. These workers are equally productive in segments C and NC, furthermore, there is a free mobility.

a) Graphically show the equilibrium W and L for both segments before and after minimum wage. Show Employer Surplus and Economic Rent before and after minimum wage.

b) Suppose we know the demand for labor in both segments: L c = 1200 - W and L nc = 1000 - 2W. Solve for competitive equilibrium W, Lc and Lnc.

c) Suppose the minimum wage is 700 in the covered segment. Solve for L c, Lnc, Wnc.

d) What is the DWL due to the minimum wage?
Answer #1

Ekonland economy is divided into two segments: C and NC, where C is covered by minimum wage, while NC is not.

firstly, we will solve B part

demand for labor in covered by minimum wage segment: Lc = 1200 - W -----------(1) and

demand for labor not covered by minimum wage segment Lnc = 1000 - 2W.--------(2)

as total number of labor = 1000

therefore, Lc + Lnc = 1000

so, Lnc = 1000 - Lc

solving above (2) equation L nc = 1000 - 2W.

1000 - 2W = 1000 - Lc

Lc = 2W

putting above value in equation 1

2W = 1200 - W

W = 400

Lc = 800

Lnc = 200

PART a graph of covered by minimum wage segment:

wage 1200 - Supply - 800 susplus Employes 4006 Tile D 1200 labor (4) Ecomowyie reut Before minimum wage Wage - 400 labor dlew

W = 400

Lc = 800

economic rent = ODBE (320000)

employer surplus = AEB( 1/2 * 800* 800) = 320000

Ates minimum wage rule 700 susplus 1200 Emplores Lunemployed =300 700 JA LC - 1200-W 16 SOO G 1200 800 D wage (new) = 700 Cab

DUE TO INCREASE IN WAGE DEMAND FOR LABOR DECREASE IN SEGMENT RESULT IN UNEMPLOYMENT OF 300 LABOR

now ,

not covered market through minimum wage market.
when minimum wage is imposed in one market demand for labor decreased in that market as we can see in above figure, due to which 300 labor unemployed now since their is mobility in labor market these 300 labor find work in other segments of market and increase labor supply by 300 .
and hence decreases wage.

before minimum wage

WAGE = 400
LABOR = 200
EMPLOYER SURPLUS = (AHB) 10000
ECONOMIC RENT = (HBFO) 80000

ware 24 Lore wase minimum B 500 400 - 200t c LNC= 100620 lol LNG (Lobos 206 so 1000 UN

AFTER MINIMUM WAGE RULE

WAGE = 250 (PUT IN LABOR DEMAND EQUATION Lnc = 500)
LABOR = 500
EMPLOYER SURPLUS = (AGC) (62500)
ECONOMIC RENT = OECG(125000)

PART. D

DEADWEIGHT LOSS IS TOTAL LOSS + GAIN IN BOTH SEGMENTS OF MARKET
IN covered through the minimum wage market. LOSS = (HGDB) = 1,65,000

IN NOT covered through the minimum wage market there IS GAIN DUE TO INCREASES IN LABOR SUPPLY =(BCEF) 97,500

SO THEIR IS DEADWEIGHT LOSS = (165000- 97500 ) = 67,500